# DGFusion: Depth-Guided Sensor Fusion for Robust Semantic Perception

### v1

@@ III-B Losses
Outlier-robust $L_{1}$ loss. Lidar returns degrade in adverse conditions, making the raw depth measurements noisy and unreliable. To address this, we compute an $L_{1}$ loss on logarithmic depth values. The $L_{1}$ loss is inherently more robust to outliers than $L_{2}$, and using the logarithmic scale better handles depth variations across orders of magnitude [5, 29, 30]. Let $\hat{D}_{p}$ be the predicted depth at pixel $p$ and $D_{p}$ the corresponding raw lidar reference depth. Since raw lidar depths are often severely corrupted in adverse conditions, directly supervising depth predictions using all available measurements can degrade learning (cf. Sec. IV-B). To address this, we restrict the loss computation to a more reliable subset of pixels, i.e. those pixels below the $\tau$-quantile of depth errors. We first define the absolute log-error at pixel $p$ as
@@
Edge-aware smoothness loss. To guide depth prediction in regions lacking reliable lidar supervision, we employ an edge-aware smoothness term popular in monocular depth estimation [26]. Specifically, we weigh depth gradients with the local intensity gradients in the RGB image, thereby encouraging smooth depth predictions while preserving discontinuities at actual object boundaries:
@@
Panoptic-edge-aware smoothness loss. In our multi-task setup, we leverage ground-truth panoptic segmentation labels to guide depth estimation. In depth maps, sharp discontinuities often coincide with object boundaries, which include both semantic class boundaries and instance-level boundaries within the same class. Previous semantic-aware smoothness losses [24] only handle class transitions, e.g. between “road” and “car”, while several scenes also contain multiple adjacent objects of the same class, e.g. two adjacent “cars”, whose shared boundaries are equally informative about depth transitions. To overcome this limitation, we extend the semantic-edge formulation to a panoptic-edge one, ensuring that depth gradients respect both class and instance boundaries. In particular, we compute a binary mask based on the panoptic labels to determine where depth gradients should be penalized. More formally, let
@@
In this formulation, $w^{x}_{p}$ and $w^{y}_{p}$ are 0 (i) around panoptic boundaries, allowing depth discontinuities, and (ii) at unlabeled regions, where variations in depth are not penalized. This promotes smooth depth estimates in the interior of each object. Thus, our loss in Eq. 8 guides the depth estimation to smoothly fill in areas between valid lidar pixels supervised via $\mathcal{L}_{\text{logL1}}$, while allowing sharp transitions between adjacent objects.
Total loss. Finally, the outlier-robust depth loss is weighted together with our smoothness losses as follows:
@@
where $\lambda_{\text{L1}}$, $\lambda_{\text{es}}$, and $\lambda_{\text{pes}}$ are hyperparameters that control the relative strength of each term. We now define our total multi-task loss as
@@
where $\mathcal{L}_{\text{seg}}$ denotes the full segmentation loss used in OneFormer [17], which supervises our predicted semantics, $\mathcal{L}_{\text{cond}}$ supervises the CT embedding following CAFuser [1], and $\mathcal{L}_{\text{depth}}$ supervises our predicted depth.
@@
Mask2Former [21] C Swin-T 49.35 39.38 48.84 46.48 45.39 45.1 31.29 58.23 78.34 58.01 46.89
MaskDINO [16] C Swin-T 51.85 42.73 54.05 46.20 46.23 48.54 38.64 57.29 80.51 60.28 49.44
OneFormer [17] C Swin-T 57.55 47.83 58.33 53.68 53.43 53.77 43.54 63.69 80.93 66.97 55.21
HRFuser [6] CLRE HRFuser-T 44.60 40.01 47.03 43.59 42.69 40.62 28.33 55.23 78.41 54.34 43.90
MUSES [11] CLRE 4xSwin-T 54.06 49.73 55.28 50.34 53.77 50.51 39.94 63.53 81.06 65.02 53.60
CAFuser-CAA [1] CLRE Swin-T 59.93 56.24 61.16 56.41 59.38 57.88 48.03 67.64 81.80 71.61 59.38
CAFuser [1] CLRE Swin-T 59.49 57.34 61.36 57.52 59.63 57.2 48.42 67.90 82.03 71.75 59.70
@@ III-C Implementation Details
We use a Swin-T backbone [31] pre-trained on ImageNet [32] and follow the training protocol of CAFuser [1]. DGFusion is trained with a batch size of 8 for 180k iterations on MUSES and 200k on DeLiVER. For the depth branch, we set $\tau=0.8$, $\lambda_{\text{L1}}=0.9$, $\lambda_{\text{es}}=0.05$, $\lambda_{\text{pes}}=0.05$, and $\lambda_{\text{depth}}=1$. We use the OneFormer head [17] for our semantic branch, which enables semantic, instance, and panoptic segmentation via a single model, and we adopt the default settings for the segmentation and CT losses from OneFormer and CAFuser, respectively. All input modalities are normalized over the entire dataset and the data augmentations from CAFuser are applied. Our network is optimized using AdamW with a poly learning rate scheduler.
@@
Mask2Former [21] C Swin-T 70.7
SegFormer [33] C MiT-B2 72.5
OneFormer [17] C Swin-T 72.8
CMNeXt [7] CLRE MiT-B2 72.1
GeminiFusion [9] CLRE MiT-B2 75.3
CAFuser-CAA [1] CLRE Swin-T 78.5
CAFuser [1] CLRE Swin-T 78.2
@@ IV Experiments
We evaluate DGFusion on two challenging multi-sensor datasets with a focus on adverse conditions in driving scenes: MUSES and DeLiVER. MUSES [11] is a real-world dataset that comprises 2500 multi-sensor scenes captured under eight environmental conditions (rain, snow, fog, and clear weather, each in day/night) and containing high-quality human-annotated panoptic and semantic ground truth. Each sample includes RGB camera, lidar, radar, and event data, all of which we project onto the RGB plane with the help of the official SDK. We perform all our ablations on this challenging real-world dataset. Due to computational constraints, the hyperparameter ablation in Tab. VI was performed on half of the training data. DeLiVER [7] is a synthetic dataset, featuring 7885 scenes recorded under adverse conditions (cloudy, foggy, night, rainy) and corner-case sensor failures (e.g. motion blur, lidar jitter) with semantic ground truth. Each scene provides RGB camera, lidar, and event camera modalities as well as a pixel-perfect depth map of the scenes. As this depth map is not a direct sensor recording and not realistic to obtain in this quality in any real-world scenario, we argue that one should treat it rather as an additional ground truth. Hence, we use the depth map as the dense supervision for the depth head of our method. As previous methods treat it as an input, we compare to such methods on both scenarios, with (CLDE) and without (CLE) the depth map as an input.
@@ IV-A Comparison with The State of The Art
DGFusion sets the new state of the art (SOTA) in multi-sensor panoptic and semantic segmentation compared to prior methods. As presented in Tab. I, we outperform previous methods on MUSES across all metrics, including a variety of adverse weather and lighting conditions. We set the new SOTA with 61.03% PQ (+1.32%) and improvements are most pronounced in adverse conditions. Compared with the previous SOTA method CAFuser [1], all adverse conditions show a substantial improvement: 1.34% PQ in Fog, 1.63% PQ in Rain and even +2.57% PQ in Snow conditions, compared to a moderate +0.8% PQ in Clear conditions. This confirms our hypothesis that fusing both global condition cues and local depth features is crucial for robust performance under adverse environmental conditions.
@@
CMNeXt [7] MiT-B2 50.3 53.0
StitchFusion [8] MiT-B2 50.8 53.4
GeminiFusion [9] MiT-B2 50.5 54.5
CAFuser-CAA [1] Swin-T 51.2 55.2
CAFuser [1] Swin-T 51.3 55.6
@@
As the OneFormer segmentation head can predict both panoptic and semantic segmentation, we also compare DGFusion to SOTA multimodal semantic segmentation methods on the MUSES dataset. Tab. II shows that our method significantly outperforms both camera-only baselines and prior multimodal approaches, achieving 79.5% mIoU (+ 1.0%). Notably, we surpass other fusion-based methods, such as CMNeXt [7] and CAFuser [1], highlighting the effectiveness of our multi-task setup for improving semantic segmentation in adverse-weather driving scenes.
@@
IV-B Ablations
We perform ablation studies to investigate the impact of each proposed architectural component in Tab. IV, supervised with the complete loss described in Sec. III-B. Starting with the baseline CAFuser (ID 1), adding an auxiliary depth head (ID 2) significantly boosts performance by +0.64% PQ. Notably, this auxiliary depth head is used only during training and does not alter the inference architecture at all, indicating that leveraging existing lidar inputs as auxiliary supervision provides a substantial benefit. When further adding our proposed depth-guided fusion module (ID 4), performance improves by an additional +0.57% PQ. We also observe a synergy of the global CT and local DT, as removing either the global CT (ID 3) or the local DT (ID 2) leads to performance drops of -0.53% PQ and -0.57% PQ, respectively. These results confirm that localized depth cues and condition-aware fusion mutually reinforce robust multimodal segmentation, demonstrating the individual contributions and complementary strengths of our proposed architectural components.
TABLE IV: Ablation of architectural components on the MUSES test set.
ID CT Aux. Depth Head DT mIoU $\uparrow$ PQ $\uparrow$
1 - - 78.20 59.70
2 - 78.63 60.34
3 - 79.19 60.38
4 (Ours) 79.50 61.03
TABLE V: Loss design ablations on the MUSES test set.
1 - - 78.87 59.84
2 - 78.94 59.98
3 - 78.50 60.04
4 (Ours) 79.50 61.03
@@ IV-B Ablations
MUSES [11]
OneFormer [17]
CAFuser [1]
@@
We also investigate our depth loss design from Sec. III-B, with results presented in Tab. V. The baseline (ID 1), which is trained only with a basic $L_{1}$ loss as depth supervision, achieves moderate performance. Introducing our smoothness terms $\mathcal{L}_{\text{es}}$ and $\mathcal{L}_{\text{pes}}$ (ID 2) improves PQ slightly by +0.14%, indicating their benefit in regularizing sparse lidar depth supervision. Finally, combining both our smoothness terms and $\tau$-quantile filtering (ID 4) results in the best overall performance, further improving PQ by +0.93%. Our smoothness losses shine especially at the presence of $\tau$-filtering, as removing the former (from ID 4 to ID 3) reduces PQ by -0.87%. These experiments confirm the importance of explicitly addressing sparse and noisy lidar returns through outlier filtering and smoothness constraints for more robust depth guidance and hence improved panoptic segmentation.
@@
Finally, in Tab. VII, we investigate the computational efficiency of our method. DGFusion adds only 2.3% to the parameter count at test time and runs at 6.83 FPS on an A6000 GPU. Hence, our significant +1.21% PQ gain comes at only marginal computational cost. This demonstrates that our depth-guided fusion approach achieves substantial performance improvements while maintaining practical efficiency for real-world deployment.
@@ IV-C Qualitative Results
In Fig. 3, we present qualitative prediction results on MUSES under various conditions, demonstrating the strengths of our approach compared to existing methods. Under the clear day scenario, we accurately capture the tram in the distance and all four pedestrians on the left sidewalk, whereas other baselines overlook some of these persons. In the second example, our model successfully segments the person on the right edge and the cyclist on the left and even detects one of the smaller signs in the background. In the foggy night scene, we are the only method that detects both the cyclist and the bicycle despite the very low visibility. In the snowy nighttime case, we correctly identify wall regions on the side of the bridge in front and detect the bicycle on the left. Finally, in the snow daytime scene, DGFusion is the only method that predicts at least one of the two pedestrians on the left sidewalk, though it misses the second person, highlighting a limitation in detecting heavily occluded objects. Our method also occasionally confuses similar classes, such as “wall” and “fence” (4th and last rows). Nevertheless, these examples highlight how our depth-guided method improves detection under all environmental conditions.

### v2

@@ III-B Losses
Outlier-robust $L_{1}$ loss. Lidar returns degrade in adverse conditions, making the raw depth measurements noisy and unreliable. To address this, we compute an $L_{1}$ loss on logarithmic depth values. The $L_{1}$ loss is inherently more robust to outliers than $L_{2}$, and using the logarithmic scale better handles depth variations across orders of magnitude [rahman2024semi, piccinelli2024unidepth, piccinelli2025unidepthv2]. Let $\hat{D}_{p}$ be the predicted depth at pixel $p$ and $D_{p}$ the corresponding raw lidar reference depth. Since raw lidar depths are often severely corrupted in adverse conditions, directly supervising depth predictions using all available measurements can degrade learning (cf. Sec. IV-B). To address this, we restrict the loss computation to a more reliable subset of pixels, i.e. those pixels below the $\tau$-quantile of depth errors. We first define the absolute log-error at pixel $p$ as
@@
Edge-aware smoothness loss. To guide depth prediction in regions lacking reliable lidar supervision, we employ an edge-aware smoothness term popular in monocular depth estimation [godard2017unsupervised]. Specifically, we weigh depth gradients with the local intensity gradients in the RGB image, thereby encouraging smooth depth predictions while preserving discontinuities at actual object boundaries:
@@
Panoptic-edge-aware smoothness loss. In our multi-task setup, we leverage ground-truth panoptic segmentation labels to guide depth estimation. In depth maps, sharp discontinuities often coincide with object boundaries, which include both semantic class boundaries and instance-level boundaries within the same class. Previous semantic-aware smoothness losses [chen2019towards] only handle class transitions, e.g. between “road” and “car”, while several scenes also contain multiple adjacent objects of the same class, e.g. two adjacent “cars”, whose shared boundaries are equally informative about depth transitions. To overcome this limitation, we extend the semantic-edge formulation to a panoptic-edge one, ensuring that depth gradients respect both class and instance boundaries. In particular, we compute a binary mask based on the panoptic labels to determine where depth gradients should be penalized. More formally, let
@@
In this formulation, $w^{x}_{p}$ and $w^{y}_{p}$ are 0 (i) around panoptic boundaries, allowing depth discontinuities, and (ii) at unlabeled regions, where variations in depth are not penalized. This promotes smooth depth estimates in object interior. Thus, our loss in (15) guides depth estimation to smoothly fill in areas between valid lidar pixels supervised via $\mathcal{L}_{\text{logL1}}$, while allowing sharp transitions between adjacent objects.
Total loss. Finally, the outlier-robust depth loss is weighted together with our smoothness losses as
@@
where $\lambda_{\text{L1}}$, $\lambda_{\text{es}}$, and $\lambda_{\text{pes}}$ are hyperparameters that control the relative strength of each term. Our total multi-task loss is
@@
where $\mathcal{L}_{\text{seg}}$ denotes the full segmentation loss used in OneFormer [jain2023oneformer], which supervises our predicted semantics, $\mathcal{L}_{\text{cond}}$ supervises the CT embedding following CAFuser [cafuser], and $\mathcal{L}_{\text{depth}}$ supervises our predicted depth.
@@
Mask2Former [cheng2022masked] C Swin-T 49.35 39.38 48.84 46.48 45.39 45.1 31.29 58.23 78.34 58.01 46.89
MaskDINO [li2023mask] C Swin-T 51.85 42.73 54.05 46.20 46.23 48.54 38.64 57.29 80.51 60.28 49.44
OneFormer [jain2023oneformer] C Swin-T 57.55 47.83 58.33 53.68 53.43 53.77 43.54 63.69 80.93 66.97 55.21
HRFuser [broedermann2023hrfuser] CLRE HRFuser-T 44.60 40.01 47.03 43.59 42.69 40.62 28.33 55.23 78.41 54.34 43.90
MUSES [MUSES] CLRE 4xSwin-T 54.06 49.73 55.28 50.34 53.77 50.51 39.94 63.53 81.06 65.02 53.60
CAFuser-CAA [cafuser] CLRE Swin-T 59.93 56.24 61.16 56.41 59.38 57.88 48.03 67.64 81.80 71.61 59.38
CAFuser [cafuser] CLRE Swin-T 59.49 57.34 61.36 57.52 59.63 57.2 48.42 67.90 82.03 71.75 59.70
@@ III-C Implementation Details
We use a Swin-T backbone [liu2021swin] pre-trained on ImageNet [deng2009imagenet] and follow the training protocol of CAFuser [cafuser]. DGFusion is trained with a batch size of 8 for 180k iterations on MUSES and 200k on DeLiVER. For the depth branch, we set $\tau=0.8$, $\lambda_{\text{L1}}=0.9$, $\lambda_{\text{es}}=0.05$, $\lambda_{\text{pes}}=0.05$, and $\lambda_{\text{depth}}=1$. We use the OneFormer head [jain2023oneformer] for our semantic branch, which enables semantic, instance, and panoptic segmentation via a single model, and we adopt the default settings for the segmentation and CT losses from OneFormer and CAFuser, respectively. All input modalities are normalized over the entire dataset and the data augmentations from CAFuser are applied. Our network is optimized using AdamW with a poly learning rate scheduler.
@@
Mask2Former [cheng2022masked] C Swin-T 70.7
SegFormer [xie2021segformer] C MiT-B2 72.5
OneFormer [jain2023oneformer] C Swin-T 72.8
CMNeXt [zhang2023delivering] CLRE MiT-B2 72.1
GeminiFusion [jia2024geminifusion] CLRE MiT-B2 75.3
CAFuser-CAA [cafuser] CLRE Swin-T 78.5
CAFuser [cafuser] CLRE Swin-T 78.2
@@ IV Experiments
We evaluate DGFusion on two challenging multi-sensor datasets with a focus on adverse conditions in driving scenes: MUSES and DeLiVER. MUSES [MUSES] is a real-world dataset that comprises 2500 multi-sensor scenes captured under eight environmental conditions (rain, snow, fog, and clear weather, each in day/night) and containing high-quality human-annotated panoptic and semantic ground truth. Each sample includes RGB camera, lidar, radar, and event data, all of which we project onto the RGB plane with the help of the official SDK. We perform all our ablations on this challenging real-world dataset. Due to computational constraints, the hyperparameter ablation in Tab. VI was performed on half of the training data. DeLiVER [zhang2023delivering] is a synthetic dataset, featuring 7885 scenes recorded under adverse conditions (cloudy, foggy, night, rainy) and corner-case sensor failures (e.g. motion blur, lidar jitter) with semantic ground truth. Each scene provides RGB camera, lidar, and event camera modalities as well as a pixel-perfect depth map of the scenes. As this depth map is not a direct sensor recording and not realistic to obtain in this quality in any real-world scenario, we argue that one should treat it rather as an additional ground truth. Hence, we use the depth map as the dense supervision for the depth head of our method. As previous methods treat it as an input, we compare to such methods on both scenarios, with (CLDE) and without (CLE) the depth map as an input.
@@ IV-A Comparison with The State of The Art
DGFusion sets the new state of the art (SOTA) in multi-sensor panoptic and semantic segmentation compared to prior methods. As presented in Tab. I, we outperform previous methods on MUSES across all metrics, including a variety of adverse weather and lighting conditions. We set the new SOTA with 61.03% PQ (+1.33%) and improvements are most pronounced in adverse conditions. Compared with the previous SOTA method CAFuser [cafuser], all adverse conditions show a substantial improvement: 1.34% PQ in Fog, 1.63% PQ in Rain and even +2.57% PQ in Snow conditions, compared to a moderate +0.8% PQ in Clear conditions. This confirms our hypothesis that fusing both global condition cues and local depth features is crucial for robust performance under adverse environmental conditions.
@@
CMNeXt [zhang2023delivering] MiT-B2 50.3 53.0
StitchFusion [li2024stitchfusion] MiT-B2 50.8 53.4
GeminiFusion [jia2024geminifusion] MiT-B2 50.5 54.5
CAFuser-CAA [cafuser] Swin-T 51.2 55.2
CAFuser [cafuser] Swin-T 51.3 55.6
@@
As the OneFormer segmentation head can predict both panoptic and semantic segmentation, we also compare DGFusion to SOTA multimodal semantic segmentation methods on the MUSES dataset. Tab. II shows that our method significantly outperforms both camera-only baselines and prior multimodal approaches, achieving 79.5% mIoU (+ 1.0%). Notably, we surpass other fusion-based methods, such as CMNeXt [zhang2023delivering] and CAFuser [cafuser], highlighting the effectiveness of our multi-task setup for improving semantic segmentation in adverse-weather driving scenes.
@@
MUSES [MUSES]
OneFormer [jain2023oneformer]
CAFuser [cafuser]
@@
IV-B Ablations
We perform ablation studies to investigate the impact of each proposed architectural component in Tab. IV, supervised with the complete loss described in Sec. III-B. Starting with the baseline CAFuser (ID 1), adding an auxiliary depth head (ID 2) significantly boosts performance by +0.64% PQ. Notably, this auxiliary depth head is used only during training and does not alter the inference architecture at all, indicating that leveraging existing lidar inputs as auxiliary supervision provides a substantial benefit. When further adding our proposed depth-guided fusion module (ID 3), performance improves by an additional +0.69% PQ. We also observe a synergy of CAFuser’s global CT and our local DT, as removing either CT (ID 4) or DT (ID 2) leads to performance drops of 0.65% PQ and 0.69% PQ, respectively. These results confirm that localized depth cues and condition awareness mutually reinforce robust multimodal segmentation, demonstrating the individual contributions and complementary strengths of our proposed architectural components.
TABLE IV: Ablation of architectural components on the MUSES test set.
ID CT Aux. Depth Head DT mIoU $\uparrow$ PQ $\uparrow$
1 - - 78.20 59.70
2 - 78.63 60.34
3 (Ours) 79.50 61.03
4 - 79.19 60.38
TABLE V: Loss design ablations on the MUSES test set.
1 - - 78.87 59.84
2 - 78.94 59.98
3 - 78.50 60.04
4 (Ours) 79.50 61.03
We also investigate our depth loss design from Sec. III-B, with results presented in Tab. V. The baseline (ID 1), which is trained only with a basic $L_{1}$ loss as depth supervision, achieves moderate performance. Introducing our smoothness terms $\mathcal{L}_{\text{es}}$ and $\mathcal{L}_{\text{pes}}$ (ID 2) improves PQ slightly by +0.14%, indicating their benefit in regularizing sparse lidar depth supervision. Finally, combining both our smoothness terms and $\tau$-quantile filtering (ID 4) results in the best overall performance, further improving PQ by +1.05%. Our smoothness losses shine especially at the presence of $\tau$-filtering, as removing the former (from ID 4 to ID 3) reduces PQ by -0.99%. These experiments confirm the importance of explicitly addressing sparse and noisy lidar returns through outlier filtering and smoothness constraints for more robust depth guidance and hence improved panoptic segmentation.
@@ IV-B Ablations
Finally, in Tab. VII, we investigate the computational efficiency of our method. DGFusion adds only 2.3% to the parameter count at test time and runs at 6.83 FPS on an A6000 GPU. Hence, our significant +1.33% PQ gain comes at only marginal computational cost. This demonstrates that our depth-guided fusion approach achieves substantial performance improvements while maintaining practical efficiency for real-world deployment.
@@ IV-C Qualitative Results
In Fig. 3, we present qualitative prediction results on MUSES under various conditions, demonstrating the strengths of our approach compared to existing methods. Under the clear day scenario, we accurately capture the tram in the distance and all four pedestrians on the left sidewalk, whereas other baselines overlook some of these persons. In the second example, our model successfully segments the person on the right edge and the cyclist on the left and even detects one of the smaller signs in the background. In the foggy night scene, we are the only method that detects both the cyclist and the bicycle despite the very low visibility. Finally, in the snow daytime scene, DGFusion is the only method that predicts at least one of the two pedestrians on the left sidewalk, though it misses the second person, highlighting a limitation in detecting heavily occluded objects. Our method also occasionally confuses similar classes, such as “wall” and “fence” (last row). Nevertheless, these examples highlight how our depth-guided method improves detection under all environmental conditions.
@@ V Conclusion
TABLE VIII: Class-wise PQ results on the MUSES test split.
Method
road
sidewalk
building
wall
fence
pole
traffic light
traffic sign
vegetation
terrain
sky
person
rider
car
truck
bus
train
motorcycle
bicycle
Mask2Former 95.12 72.08 76.59 38.35 28.96 39.85 37.54 48.73 69.73 49.60 83.95 34.03 19.31 55.70 28.59 42.58 38.21 14.88 17.06
MaskDINO 95.22 69.06 77.87 31.32 32.81 44.31 37.52 44.56 70.81 42.23 84.51 46.88 27.65 62.22 35.41 41.41 38.07 30.89 26.61
OneFormer 95.95 75.68 81.29 44.13 37.30 47.80 48.21 56.87 72.37 52.53 88.50 47.07 29.92 64.41 40.94 62.03 47.37 31.14 25.44
HRFuser 94.35 69.15 71.73 34.71 24.43 35.61 30.04 51.23 67.62 46.49 82.16 33.50 18.89 56.49 26.87 31.53 29.73 19.94 9.70
MUSES 95.86 76.42 80.40 47.43 37.56 45.57 42.83 60.23 75.01 53.83 83.69 42.21 33.51 63.65 40.01 45.09 42.65 28.52 23.88
CAFuser-CAA 96.14 78.29 83.12 51.40 42.48 49.35 53.79 64.07 78.65 56.09 90.65 51.79 39.53 66.33 46.88 58.99 52.51 39.07 29.18
CAFuser 96.36 79.52 84.38 50.87 43.55 49.84 50.82 63.18 79.40 56.90 92.11 50.51 39.05 65.69 45.95 62.23 54.34 40.63 28.96
DGFusion (Ours) 96.63 80.80 84.51 53.60 43.56 52.54 54.39 65.56 79.41 57.22 93.63 51.74 42.17 67.13 49.33 64.07 53.26 39.41 28.36
Detailed Class-Level Results
We present detailed class-level panoptic quality results on the MUSES test split in Table VIII. Our method consistently achieves superior performance across most classes. This highlights our model’s robust adaptation across varied semantic categories compared to other strong multimodal fusion methods, reflecting the effectiveness of our depth-guided fusion strategy.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
RGB
Lidar
ID 1 of Tab. 5 ($\lambda_{\text{es}}=0;\lambda_{\text{pes}}=0;\tau=1$)
DGFusion-depth (Ours)
Figure 4: Visual study Qualitative ablation on our loss design on the MUSES dataset. Best viewed on screen at full zoom.
Visual Ablation on Our Loss Design
In  Figure 4, we visually explore the effect of our loss design, which we investigate quantitatively in the main paper. The results illustrate that our proposed loss enhances the depth estimation quality by targeting the challenges of using noisy lidar as ground truth. Specifically, structural artifacts such as holes in building facades are effectively reduced, and boundaries are notably cleaner. Thereby demonstrating the strength of our method in handling challenging visual scenarios and providing visually superior depth predictions.
Further Qualitative Results
In Figure 5, we present further qualitative comparisons across diverse scenarios on the MUSES dataset. Our proposed method DGFusion shows improved robustness in scenarios challenging for RGB-only methods and baseline fusion approaches, such as perception at night, fog, and detection of distant objects. Specifically, it achieves superior instance separation and effectively leverages multimodal information under adverse conditions, thereby visually confirming the quantitative improvements in panoptic segmentation accuracy.
RGB
Lidar
Radar
Events
Ours-depth
GT
MUSES
OneFormer
CAFuser
DGFusion (Ours)
Figure 5: Further qualitative results on MUSES. Best viewed on a screen at full zoom.